OpenStudy (anonymous):

solve for x 1/2+4/2x=x+4/10

OpenStudy (anonymous):

\[\left(\begin{matrix}4 \\ 2\end{matrix}\right)x\] or \[\left(\begin{matrix}4 \\ 2x\end{matrix}\right)\] ?

OpenStudy (anonymous):

thats not one of the choices

OpenStudy (anonymous):

do u mean x(1/2)+x(4/2)=x+4/10?

OpenStudy (anonymous):

If that was the question then the ans is x=1/6

OpenStudy (anonymous):

\[x = 1/20 (1-3\sqrt{89} )\] and \[x = 1/20 (1-3\sqrt{89} )\]